As a fraction 25 mins to 1.5 hours
1
Expert's answer
2020-02-05T09:37:47-0500

"1\ hour = 60\ minutes \newline 
1.5 \ hours = (1.5\times 60)\ minutes = 90\ minutes \newline fraction \ of \ 25\ minutes \ to \ 1.5\ hours\ is \ equivalent\ to \newline \dfrac{25}{90} \ = \ \dfrac{5}{18}"